1. Ball thrown at 30° with intial velocity of 20m/s.



Find the



A. Initial velocity



B. Final velocity



C. Acceleration (m/s²)



D. Time ( s )



E. Distance ( m )

1
Expert's answer
2022-01-18T13:42:33-0500

"\text{ A.}\space v_i=20\text{ m/s}.\\
\text{ B.}\space v_f=20\text{ m/s}.\\
\text{ C.}\space g=9.8\text{ m/s}^2.\\
\text{ D.}\space t=2v\sin\theta/g=2\text{ s}.\\
\text{ E.}\space D=v\cos\theta·t=34.6\text{ m}.\\"